# Epistemic Wrapping for Uncertainty Quantification

### v1

@@ 1 Introduction
This paper proposes Epistemic Wrapper, a novel method which, for the first time, models EU in the parameter space via a random set representation by “wrapping" a learnt Bayesian posterior there in the form of a belief function. posterior (Fig. 1). Our Epistemic Wrapper follows a structured five-step process, where each step is executed in a hierarchical manner: (i) We begin by extracting posterior distributions with parameters ($\mu,\sigma$) from a pre-trained BNN model, where the priors are modeled as Gaussian distributions. (ii) The posterior distributions are then truncated, and continuous belief functions are computed over closed intervals. (iii) In the third step, these belief values are transformed into mass values using Moebius inversion. (iv) A Dirichlet distribution is fitted to the grid of mass values using the Method of L-moments. (v) Finally, inference is performed using a Hybrid Interval Neural Network (INN).
Our approach leverages the strengths of BNNs while injecting the ability of higher-order measure to improve robustness and uncertainty estimation.
@@ 2.1 Bayesian Neural Networks
While various types of uncertainty measures cuzzolin2021big have been employed in machine learning in the past cuzzolin2013if ; cuzzolin18belief-maxent ; liu2019evidence ; gong2017belief , recent advancements in epistemic uncertainty modelling have introduced a range of methods to improve predictive reliability across various neural architectures. Evidential deep learning predicts second-order probability distributions to estimate uncertainty, but faces challenges in optimisation and interpretation (juergens2024is, ). Methods like G-$\Delta$UQ refine uncertainty calibration in Graph Neural Networks (GNNs) through stochastic data centering (trivedi2024accurate, ), while SPDE-based GNNs employ $Q$-Wiener processes for uncertainty propagation in complex graphs (lin2024graph, ). The Graph Energy-Based Model (GEBM) leverages graph diffusion to quantify uncertainty at different structural levels (fuchsgruber2024energybased, ), and credal set-based ensemble learning constructs plausible probability distributions to measure aleatoric and epistemic uncertainty (hofman2024quantifying, ).
@@ 3.2 Dynamic Truncation
Figure 2: Posterior distribution of weights for the last layer. The histogram displays the sampled weights, overlaid with a fitted normal distribution (mean = 0.01, std = 0.23). Vertical dashed lines indicate truncation points at ±$\pm 3\sigma$.
@@
[…]
@@ 3.3 Continuous Belief Functions on Closed Intervals
Continuous belief functions on intervals. Belief functions can be easily extended to continuous spaces (e.g., a network’s parameter space) by defining a continuous mass function over the collection of closed intervals, rather than the entire power set. Given a network parameter $\omega$ with values in $\mathbb{R}$, this requires defining a continuous PDF over the collection of intervals $[a,b]\subset\mathbb{R}$ (cuzzolin2020geometry, ). Here we will assume that parameter values are bounded after truncation (for illustration, in $[0,1]$); however, the method can be easily extended to unbounded parameter values as well.
@@
$\vspace{-6pt}Pl_{\Theta}(A|\mathbb{D})=\sup_{\omega\in A}\hat{p}(\boldsymbol{\omega}|\mathbb{D}).\vspace{-2pt}$ (3)
@@
$\vspace{-6pt}Bel_{\Theta}(A|\mathbb{D})=1-Pl_{\Theta}(A^{c}|\mathbb{D}),\vspace{-2pt}$ (4)
@@ 3.4 Fitting a Dirichlet Distribution
$\vspace{-8pt}f(x_{1},\ldots,x_{K};\alpha_{1},\ldots,\alpha_{K})=\frac{1}{B(\alpha)}\prod_{i=1}^{K}x_{i}^{\alpha_{i}-1}.\vspace{-1pt}$ (5)
@@
0:  $\mathbf{X}\in\mathbb{R}^{n\times 3}$ (3D simplex coordinates), $\mathbf{w}\in\mathbb{R}^{n}$ (masses), $\epsilon>0$
@@ 3.6 Inference via Hybrid Interval Neural Networks
Namely, the unwrapped weights (Gaussian posteriors) generate the following intervals: $\text{Lower bound}=\mu-\sigma,\text{Upper bound}=\mu+\sigma.$ The process of defining these wrapped and unwrapped weights can be considered the weight initialisation step for the Hybrid-INN model. In contrast, the baseline Hybrid-INN retains the default weight initialisation scheme (kim1993understanding, ). In other words, at inference time, the model operates as a Hybrid-INN but with two distinct sets of weights: (i) with random initialisation, (baseline Hybrid-INN) and (ii) with weights transformed by the Epi-Wrapper. Both models are also fine-tuned: the baseline with randomly initialized weights, while the Epi-Wrapper utilizes weights transformed through the wrapping strategy. For a fair comparison, both the baseline and our Epi-Wrapper operate under identical functional settings.
@@ 4 Experiments
Hybrid-INN Epi-Wrapper BNN Hybrid-INN Epi-Wrapper
@@ 4.2 Ablation on Budgeting
Budgeting using High Mean ($\uparrow\mu$) is another strategy in which we sample and “wrap" the $5\%$ weights with ’High Mean’ from the posterior distributions. From the results shown in Table 1, it can be seen that “High Mean" performs better for MLP size (no hidden units) $=8$.
[…]
@@
Hybrid-INN Epi-Wrapper BNN Hybrid-INN Epi-Wrapper
@@ 4.3 Fine-Tuning
We performed the Fine-tuning of the models, the baseline (Hybrid-INN) and ours (Epi-Wrapper), on the training data. The results are shown in Table 1. In comparison to the Hybrid-INN and BNN baselines, our model performs well as the weights wrapping mechanism acts as an initialization strategy in fine-tuning.
@@
iD Test Accuracy (%) AUROC AUPRC EU
Hybrid-INN 91.07 $\pm$ 0.08 0.5329 0.8951 0.0023
@@
Analysis of Accuracy Rejection Curve (ARC) on in-Domain Data Figure 6 presents the Accuracy Rejection Curve (ARC) for the in-domain (iD) dataset, comparing the performance of two models, Hybrid-INN and Epi-Wrapper, on MNIST. In the in-domain setting, both the training and testing are conducted on the MNIST dataset, with no distributional shift. It can be seen in the plot that, as the rejection rate increases, the accuracy steadily improves for both models, demonstrating that removing high-uncertainty predictions enhances the overall correctness of classification. However, the Epi-Wrapper model consistently outperforms the Hybrid-INN model across the entire range of rejection rates. This suggests that Epi-Wrapper is better at identifying and filtering uncertain predictions, leading to an increase in accuracy. Since this evaluation is conducted in an in-domain setting (iD), the observed performance improvements primarily reflect how well the models handle uncertainty within a familiar data distribution. For OOD samples, the results are presented in Table 3. Clearly our Epi-wrapper performs well in comparison to the baseline (Hybrid-INN).
Figure 6: Accuracy Rejection Curve (ARC) for the MNIST dataset, comparing Hybrid-INN and Epi-Wrapper.
@@
Table 3 presents a comparative analysis of uncertainty estimation performance between Hybrid-INN and Epi-Wrapper on out-of-distribution (OoD) samples from the Fashion-MNIST dataset, while both models were trained and evaluated in-domain (iD) on MNIST. The table provides key evaluation metrics, including iD test accuracy, AUROC (Area Under the Receiver Operating Characteristic Curve), AUPRC (Area Under the Precision-Recall Curve), and Epistemic Uncertainty. For iD Test Accuracy: Both models achieve comparable in-domain classification accuracy on MNIST, with Hybrid-INN attaining an accuracy of 91.07% and Epi-Wrapper slightly outperforming it with 91.12%. AUROC on OoD Samples: This is a key metric for OoD detection, measuring the model’s ability to distinguish between iD and OoD samples. A higher AUROC indicates better separation. Epi-Wrapper achieves a higher AUROC of 0.6673, outperforming Hybrid-INN (0.5329). This suggests that Epi-Wrapper is more effective at recognizing and distinguishing OoD samples from in-domain data. AUPRC on OoD Samples: AUPRC provides insight into how well the model prioritizes high-confidence predictions for OoD detection. Epi-Wrapper achieves an AUPRC of 0.9120, surpassing Hybrid-INN (0.8951). The higher AUPRC indicates that Epi-Wrapper generates better-calibrated uncertainty estimates, ensuring that truly OoD samples receive higher uncertainty scores.
EU represents the model’s epistemic uncertainty measure, where higher values suggest better sensitivity to OoD data. Epi-Wrapper achieves an EU of 0.0059, which is more than twice the EU of Hybrid-INN (0.0023). This implies that Epi-Wrapper assigns greater uncertainty to OoD samples, making it more reliable in real-world scenarios where detecting unfamiliar inputs is crucial.
@@ 4.5 Comparative Analysis
Table 2 offers a comparative analysis of classification accuracies on the MNIST and Fashion-MNIST datasets. Epi-Wrapper exhibits distinct performance advantages over its counterparts, BNN and Hybrid-INN. Before fine-tuning, Epi-Wrapper significantly outperforms Hybrid-INN on both datasets, achieving high accuracy of 51.33% on MNIST and 26.93% on Fashion-MNIST. This indicates our Wrapper’s superior ability to perform better on challenging data. After the fine-tuning process, while the BNN and Hybrid-INN models show notable improvements demonstrating their adaptability through training, the Epi-Wrapper maintains competitive, high-performance levels. Its superiority on the Bayesian baseline, both before and after fine tuning, confirms our hypothesis that employing higher-order, random-set representations in the parameter space is advantageous.
@@ A.1 Interval Neural Networks (INNs)
where $\oplus$, $\ominus$, and $\odot$ represent interval addition, subtraction, and multiplication, respectively [18]. The terms $[\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l}$, $[\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l-1}$, $[\underline{\boldsymbol{\omega}},\overline{\boldsymbol{\omega}}]^{l}$, and $[\underline{\boldsymbol{b}},\overline{\boldsymbol{b}}]^{l}$ denote the interval-formed outputs of the $l^{th}$ and $(l-1)^{th}$ layers, as well as the intervals of weights and biases of the $l^{th}$ layer, respectively. ${\sigma^{l}(\cdot)}$ is the activation function of the $l^{th}$ layer, which must be monotonically increasing. The application of interval arithmetic [18] in (9) grants INNs the ‘set constraint’ property. Specifically, for any $\boldsymbol{a}^{l-1}\!\in\![\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l-1}$, $\boldsymbol{\omega}^{l}\!\in\![\underline{\boldsymbol{\omega}},\overline{\boldsymbol{\omega}}]^{l}$, and $\boldsymbol{b}^{l}\!\in\![\underline{\boldsymbol{b}},\overline{\boldsymbol{b}}]^{l}$, the constraint in (10) consistently holds.
@@
If $[\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]$ is non-negative, such as the output of RELU activation, the calculation of $[\underline{\boldsymbol{o}},\overline{\boldsymbol{o}}]$ in (9) can be simplified as:
$\underline{\boldsymbol{o}}&\!=\!\text{min}\{\underline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\overline{\boldsymbol{a}}+\text{max}\{\underline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\underline{\boldsymbol{a}}\\ \overline{\boldsymbol{o}}&\!=\!\text{max}\{\overline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\overline{\boldsymbol{a}}+\text{min}\{\overline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\underline{\boldsymbol{a}}\end{aligned}.$ (11)
@@ A.3 Bayesian Baselines
For baselines, we utilize two standard variational BNNs: BNNR (Auto-Encoding Variational Bayes [26] with the local re-parameterization trick [39]), and BNNF (Flipout gradient estimator with the negative evidence lower bound loss [54]). The results before fine-tuning are presented in Table 6.
@@ A.5 Hyperparameter Analysis
Table 6: Classification accuracies for Hybrid-INN and Epi-Wrapper across CIFAR-10 and CIFAR-100 datasets using different Bayesian backbones (BNNF and BNNR) and network architectures (LeNet-5, ResNet-18, VGG-16).
Dataset Backbone # Parameters BNN Baseline Hybrid-INN Epi-Wrapper
CIFAR-10 LeNet-5 166.3K BNNF 11.01 ± 0.40% 45.54 ± 0.03%
BNNR
ResNet-18 9.82M BNNF 10.00 ± 0.10% 37.66 ± 1.10%
BNNR 9.57 ± 0.20% 35.66 ± 1.09%
VGG-16 30.24M BNNF 9.98 ± 0.21% 30.12 ± 0.78%
BNNR 9.89 ± 0.60% 31.04 ± 0.69%
CIFAR-100 ResNet-18 9.8M BNNF 10.42 ± 0.44% 21.09 ± 0.14%
BNNR 10.08 ± 0.34% 22.87 ± 0.11%
VGG-16 30.24M BNNF 9.01 ± 0.56% 19.98 ± 1.30%
BNNR 10.01 ± 0.32% 20.10 ± 1.21%

### v2

@@ 1 Introduction
This paper proposes Epistemic Wrapper, a novel method which, for the first time, models EU in the parameter space via a random set representation by “wrapping" a learnt Bayesian posterior there in the form of a belief function. posterior (Fig. 1). Our Epistemic Wrapper follows a structured five-step process, where each step is executed in a hierarchical manner: (i) We begin by extracting posterior distributions with parameters ($\mu,\sigma$) from a pre-trained BNN model, where the priors are modeled as Gaussian distributions. (ii) The posterior distributions are then truncated, and continuous belief functions are computed over closed intervals. (iii) In the third step, these belief values are transformed into mass values using Moebius inversion. (iv) A Dirichlet distribution is fitted to the grid of mass values using the Method of L-moments. (v) Finally, inference is performed using a Hybrid Interval Neural Network (INN). Our approach leverages the strengths of BNNs while injecting the ability of higher-order measure to improve robustness and uncertainty estimation.
@@ 2.1 Bayesian Neural Networks
[…]
@@ 3.2 Dynamic Truncation
Figure 2: Posterior distribution of weights for the last layer. The histogram displays the sampled weights, overlaid with a fitted normal distribution (mean = 0.01, std = 0.23). Vertical dashed lines indicate truncation points at $\pm 3\sigma$.
@@
The truncation bounds are calculated as: $\text{Lower Bound}=\mu-\text{dynamic\_multiplier}\cdot\sigma,\quad\text{Upper Bound}=\mu+\text{dynamic\_multiplier}\cdot\sigma,$ where $\mu$ is the mean, $\sigma$ is the standard deviation, and the dynamic_multiplier is calculated as: $\text{dynamic\_multiplier}=\min(5.0,\frac{1.0}{\sigma}),$ ensuring that the multiplier decreases for low-variance distributions while capping its value at 5.0 to prevent excessive truncation in high-variance cases. We have selected this approach as it provides a balance between capturing the significant probability mass of the distribution and avoiding overly wide or narrow bounds, which could either dilute meaningful mass representation or exclude critical probabilistic regions.
@@ 3.3 Continuous Belief Functions on Closed Intervals
Continuous belief functions on intervals. Belief functions can be easily extended to continuous spaces (e.g., a network’s parameter space) by defining a continuous mass function over the collection of closed intervals, rather than the entire power set. Given a network parameter $\omega$ with values in $\mathbb{R}$, this requires defining a continuous PDF over the collection of intervals $a,b]\subset\mathbb{R}$ (cuzzolin2020geometry, ). Here we will assume that parameter values are bounded after truncation (for illustration, in $[0,1]$); however, the method can be easily extended to unbounded parameter values as well.
@@
$\vskip-6.0ptPl_{\Theta}(A|\mathbb{D})=\sup_{\omega\in A}\hat{p}(\boldsymbol{\omega}|\mathbb{D}).\vskip-2.0pt$ (3)
@@
$\vskip-6.0ptBel_{\Theta}(A|\mathbb{D})=1-Pl_{\Theta}(A^{c}|\mathbb{D}),\vskip-2.0pt$ (4)
@@ 3.4 Fitting a Dirichlet Distribution
$\vskip-8.0ptf(x_{1},\ldots,x_{K};\alpha_{1},\ldots,\alpha_{K})=\frac{1}{B(\alpha)}\prod_{i=1}^{K}x_{i}^{\alpha_{i}-1}.\vskip-1.0pt$ (5)
@@
[…]
@@ 3.6 Inference via Hybrid Interval Neural Networks
Namely, the unwrapped weights (Gaussian posteriors) generate the following intervals: $\text{Lower bound}=\mu-\sigma,\text{Upper bound}=\mu+\sigma.$ The process of defining these wrapped and unwrapped weights can be considered the weight initialisation step for the Hybrid-INN model. In contrast, the baseline INN retains the default weight initialisation scheme (kim1993understanding, ). In other words, at inference time, the model operates as a Hybrid-INN but with two distinct sets of weights: (i) with random initialisation, (baseline INN) and (ii) with weights transformed by the Epi-Wrapper. Both models are also fine-tuned: the baseline with randomly initialized weights, while the Epi-Wrapper utilizes weights transformed through the wrapping strategy. For a fair comparison, both the baseline and our Epi-Wrapper operate under identical functional settings.
@@ 4 Experiments
INN Epi-Wrapper BNN INN Epi-Wrapper
@@ 4.2 Ablation on Budgeting
Budgeting using High Mean ($\uparrow\mu$) is another strategy in which we sample and “wrap" the $5\%$ weights with ’High Mean’ from the posterior distributions. From the results shown in Table 1, it can be seen that “High Mean" performs better for MLP size (no hidden units) =8=8.
In Budgeting using High Mean and High Variance ($\uparrow(\mu,\sigma$)) we rank the parameters by computing a combined score, defined as the sum of the mean and variance of their posterior distributions: $\text{combined\_score}=\mu+\sigma.$ This acts as a proxy for an upper bound of the posterior distribution, allowing us to prioritize parameters that are either highly informative (high mean) or uncertain (high variance). We then wrap these top 5% weights using Epi-wrapper. The results are shown in the Table 1. This strategy allows us to selectively wrap the most influential and uncertain parameters, ensuring that the transformation captures meaningful epistemic uncertainty. However, this approach also imposes a strict constraint on the selection process, as only weights satisfying both conditions are chosen, which may limit flexibility in certain scenarios.
@@
INN Epi-Wrapper BNN INN Epi-Wrapper
@@ 4.3 Fine-Tuning
We performed the Fine-tuning of the models, the baseline (INN) and ours (Epi-Wrapper), on the training data. The results are shown in Table 1. In comparison to the INN and BNN baselines, our model performs well as the weights wrapping mechanism acts as an initialization strategy in fine-tuning.
@@
iD Test
Accuracy (%)
AUROC AUPRC EU
INN 91.07 $\pm$ 0.08 0.5329 0.8951 0.0023
@@
Analysis of Accuracy Rejection Curve (ARC) on in-Domain Data Figure 6 presents the Accuracy Rejection Curve (ARC) for the in-domain (iD) dataset, comparing the performance of two models, INN and Epi-Wrapper, on MNIST. In the in-domain setting, both the training and testing are conducted on the MNIST dataset, with no distributional shift. It can be seen in the plot that, as the rejection rate increases, the accuracy steadily improves for both models, demonstrating that removing high-uncertainty predictions enhances the overall correctness of classification. However, the Epi-Wrapper model consistently outperforms the INN model across the entire range of rejection rates. This suggests that Epi-Wrapper is better at identifying and filtering uncertain predictions, leading to an increase in accuracy. Since this evaluation is conducted in an in-domain setting (iD), the observed performance improvements primarily reflect how well the models handle uncertainty within a familiar data distribution. For OOD samples, the results are presented in Table 3. Clearly our Epi-wrapper performs well in comparison to the baseline (INN).
Figure 6: Accuracy Rejection Curve (ARC) for the MNIST dataset, comparing INN and Epi-Wrapper.
@@
Table 3 presents a comparative analysis of uncertainty estimation performance between INN and Epi-Wrapper on out-of-distribution (OoD) samples from the Fashion-MNIST dataset, while both models were trained and evaluated in-domain (iD) on MNIST. The table provides key evaluation metrics, including iD test accuracy, AUROC (Area Under the Receiver Operating Characteristic Curve), AUPRC (Area Under the Precision-Recall Curve), and Epistemic Uncertainty. For iD Test Accuracy: Both models achieve comparable in-domain classification accuracy on MNIST, with INN attaining an accuracy of 91.07% and Epi-Wrapper slightly outperforming it with 91.12%. AUROC on OoD Samples: This is a key metric for OoD detection, measuring the model’s ability to distinguish between iD and OoD samples. A higher AUROC indicates better separation. Epi-Wrapper achieves a higher AUROC of 0.6673, outperforming INN (0.5329). This suggests that Epi-Wrapper is more effective at recognizing and distinguishing OoD samples from in-domain data. AUPRC on OoD Samples: AUPRC provides insight into how well the model prioritizes high-confidence predictions for OoD detection. Epi-Wrapper achieves an AUPRC of 0.9120, surpassing INN (0.8951). The higher AUPRC indicates that Epi-Wrapper generates better-calibrated uncertainty estimates, ensuring that truly OoD samples receive higher uncertainty scores.
EU represents the model’s epistemic uncertainty measure, where higher values suggest better sensitivity to OoD data. Epi-Wrapper achieves an EU of 0.0059, which is more than twice the EU of INN (0.0023). This implies that Epi-Wrapper assigns greater uncertainty to OoD samples, making it more reliable in real-world scenarios where detecting unfamiliar inputs is crucial.
@@ 4.5 Comparative Analysis
Table 2 offers a comparative analysis of classification accuracies on the MNIST and Fashion-MNIST datasets. Epi-Wrapper exhibits distinct performance advantages over its counterparts, BNN and INN. Before fine-tuning, Epi-Wrapper significantly outperforms INN on both datasets, achieving high accuracy of 51.33% on MNIST and 26.93% on Fashion-MNIST. This indicates our Wrapper’s superior ability to perform better on challenging data. After the fine-tuning process, while the BNN and INN models show notable improvements demonstrating their adaptability through training, the Epi-Wrapper maintains competitive, high-performance levels. Its superiority on the Bayesian baseline, both before and after fine tuning, confirms our hypothesis that employing higher-order, random-set representations in the parameter space is advantageous.
@@ A.1 Interval Neural Networks (INNs)
where $\oplus$, $\ominus$, and $\odot$ represent interval addition, subtraction, and multiplication, respectively [18]. The terms $\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l}$, $\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l-1}$, $\underline{\boldsymbol{\omega}},\overline{\boldsymbol{\omega}}]^{l}$, and $\underline{\boldsymbol{b}},\overline{\boldsymbol{b}}]^{l}$ denote the interval-formed outputs of the $l^{th}$ and $(l-1)^{th}$ layers, as well as the intervals of weights and biases of the $l^{th}$ layer, respectively. $\sigma^{l}(\cdot)$} is the activation function of the $l^{th}$ layer, which must be monotonically increasing. The application of interval arithmetic [18] in (9) grants INNs the ‘set constraint’ property. Specifically, for any $\boldsymbol{a}^{l-1}\!\in\![\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]^{l-1}$, $\boldsymbol{\omega}^{l}\!\in\![\underline{\boldsymbol{\omega}},\overline{\boldsymbol{\omega}}]^{l}$, and $\boldsymbol{b}^{l}\!\in\![\underline{\boldsymbol{b}},\overline{\boldsymbol{b}}]^{l}$, the constraint in (10) consistently holds.
@@
If $\underline{\boldsymbol{a}},\overline{\boldsymbol{a}}]$ is non-negative, such as the output of RELU activation, the calculation of $\underline{\boldsymbol{o}},\overline{\boldsymbol{o}}]$ in (9) can be simplified as:
$\begin{aligned} \underline{\boldsymbol{o}}&\!=\!\text{min}\{\underline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\overline{\boldsymbol{a}}+\text{max}\{\underline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\underline{\boldsymbol{a}}\\ \overline{\boldsymbol{o}}&\!=\!\text{max}\{\overline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\overline{\boldsymbol{a}}+\text{min}\{\overline{\boldsymbol{\omega}},\boldsymbol{0}\}\!\cdot\!\underline{\boldsymbol{a}}\end{aligned}.$ (11)
@@ A.3 Bayesian Baselines
For baselines, we utilize two standard variational BNNs: BNNR (Auto-Encoding Variational Bayes [26] with the local re-parameterization trick [39]), and BNNF (Flipout gradient estimator with the negative evidence lower bound loss [54]). The results before fine-tuning are presented in Table LABEL:table:full_cifar10_cifar100_results.